A Complete Bibliography of Internet Engineering Note (IEN) Documents

Nelson H. F. Beebe Department of Mathematics University of Utah Salt Lake City, UT 84112 USA Tel: +1 801 581 5254 FAX: +1 801 581 4148 E-mail: beebe@math.utah.edu (Internet) 12 August 2013 Version 1.03

Title word cross-reference

1 [22, 33]. 10 [121]. 11/23/34/35/40 [136]. 13 [66]. 14 [65, 145]. 15 [3, 68]. 18 [69]. 1977 [3, 65, 66]. 1978 [22, 33, 53, 63, 67, 68, 69, 70]. 1979 [76, 77, 106, 121]. 1980 [134, 145, 160]. 1981 [175]. 2 [53]. 20 [176]. 25&26 [76]. 28 [175]. 28-29-30 [175]. 29 [77]. 30 [63, 67, 175]. 4 [55, 70, 130, 134]. 40 [136]. 7 [160]. 7-8-9 [160]. 8 [106]. 9 [160]. Above [153, 154]. Access [58, 190]. Accessing [187]. Action [177].

ADA079730 [129]. Address [82, 91, 115]. Addresses [23, 31]. Addressing [16, 19, 46, 62, 83, 110, 122, 135, 162, 165, 170, 171, 174, 178, 179, 183, 188, 212]. Application [84]. Area [60, 212, 78]. Arithmetic [74]. ARPA [105, 126, 135]. ARPA/ETHERNET [126]. ARPANET [62, 171]. Assessment [99]. Assigned [93, 117, 127]. August [3, 53]. AUTODIN [101, 102].

Based [169, 193]. Bay [78]. Beyond [11]. Bibliography [37]. Book [153, 154, 155]. Boston [60]. Broadcast [10]. Buffer [182]. Build [109]. bus [130].

Cable [130]. Cable-bus [130]. Cablenet [96]. Calculation [56]. Calling [150]. Campus [143, 144]. Campus-Wide [143, 144]. Catenet [32, 48, 51, 105, 156]. CCITT [47]. Center [103]. Checksum [45, 49, 56]. Circuits [43]. CMCC [132, 157]. CMOS [164]. Collection [14]. College [185]. Comments [2, 49, 79, 177, 191, 195]. Communication [4]. Communications [207]. Community [126]. Comparative [50]. Comparison [100, 102, 130, 138]. Component [32]. Computation [159]. Conferencing [4]. Congestion [87]. Connection [8, 193]. Considerations [143]. Contents [94, 118]. Control [14, 32, 40, 55, 58, 81, 87, 105, 112, 124, 129, 190]. Controlled [156]. Controller [34]. CRC [56].

DARPA [208]. Data [14, 17, 39, 47, 207]. Datagram [1, 71, 80, 97]. Datagrams [17, 43]. Datgram [88]. DCEC [186]. DCNET [173, 194]. December [70]. DECSYSTEM [176]. DECSYSTEM-20 [176]. Delay [163]. Description [168]. Design [45, 166, 167, 9]. Development [57, 185, 198]. DIN [100]. Discuss [60]. Discussion [58, 122]. Distributed [159]. Document [167]. Dod [207, 152, 123, 124, 128, 129]. Draft [28, 47]. DTE [174]. Dynamic [25].

Early [159]. Echo [163]. EGP [209]. emption [125]. Encapsulation [140]. Enriched [62]. Environment [110, 135, 138, 143]. EPRs [138]. ETHERNET [126]. Evaluation [50]. Expanders [83]. Experience [159]. Experimental [103]. Experiments [8, 78]. Expressways [147]. Extended [86, 198]. Extensible [16]. Extensions [136]. Exterior [209].

Facilities [14]. Facility [79]. Fault [104]. February [22, 134]. Field [16]. File [36, 59, 149]. Final [151]. Flexible [97]. Flow [87]. Flying [146]. Formal [209]. Format [26, 27, 82]. Formats [44, 157, 158]. Forwarding [180]. Fragmentation [20]. FTP [99, 100, 101, 102]. Fuel [122]. Function [45].

Gateway [8, 24, 25, 30, 32, 109, 131, 181, 208, 209]. Gateways [6, 60, 196].

General [4]. Generator [172]. GGP [163]. GNOME [13, 34]. Goals [201]. Group [60]. Guide [13].

Handbook [94, 118]. Header [26, 27, 44]. Holy [137]. Host [9, 12, 192, 197, 198, 207]. Host-to-Host [12, 207]. Host/Satnet [192]. Host/SIMP [9]. Hosts [135, 173, 139]. HP3000 [167].

IEN [154]. II [31]. Implementation [30, 59, 73, 98, 168]. IMPs [139]. Informal [58]. Information [42, 103]. Initial [138]. Inter [19, 20]. Inter-Network [19, 20]. Interactions [60]. Interconnection [1, 43]. Interface [176]. Interfaces [6]. Interim [62]. Internet [2, 3, 10, 22, 26, 33, 46, 53, 60, 61, 62, 63, 76, 80, 85, 86, 87, 89, 94, 106, 108, 110, 111, 113, 118, 119, 120, 121, 123, 128, 134, 135, 138, 144, 145, 147, 158, 160, 175, 180, 184, 187, 200, 201, 208]. Internetting [11, 184, 187, 188, 189]. Internetwork [28, 40, 41, 54, 55, 78, 140]. Internetworking [48]. Involving [196]. IP [166, 170, 176, 186, 212]. IP-Addressing [170]. IP-Local [212]. IPRs [138]. ISI [211]. Isolation [104]. Issue [52, 73]. issued [199]. Issues [1, 10, 12, 122, 141, 181, 182, 184, 187, 188, 189, 196, 212]. Items [177].

January [67, 76, 77, 175, 177]. JSYS [150]. July [65]. June [68].

Latest [44]. LCS [82]. Local [212]. Logical [183]. London [185].

Mail [169, 194]. Management [136, 182, 193]. Mappings [91, 115]. March [64]. May [33, 106, 145]. Measurement [157, 161]. Measurements [163]. Mechanism [75]. Mechanisms [193]. Meeting [3, 22, 33, 53, 60, 63, 64, 65, 66, 67, 68, 69, 70, 76, 77, 104, 106, 121, 126, 134, 145, 160, 175, 177]. Memory [136, 198]. Memos [191]. Message [85, 113, 141, 157, 172, 180]. Micro [136]. Minutes [104]. MITRE [96]. Mobile [135]. Model [32, 48]. Modelling [184]. Modern [84]. Monitoring [32, 105, 131, 197]. MOS [198]. MOSIS [180]. Multi [14, 178]. Multi-net [14]. Multi-Network [178]. Multiplexing [52, 72, 90]. Mutual [140].

Name [61, 89, 103, 116]. Names [23, 31]. Naming [19, 110]. NBS [195]. NCP [11]. Net [82, 14]. Nets [147]. Network [4, 6, 19, 20, 78, 103, 120, 130, 146, 174, 178, 185, 212]. Networking [168]. Networks [1, 47, 52, 143, 207]. Never [199]. NI [99, 100, 102]. NICNAME [103]. NIFTP [169]. NIFTP-Based [169]. No [129]. Non [196]. Non-Routing [196]. Note [29]. Notes [3, 22, 33, 53, 63, 64, 65, 66, 67, 68, 69, 70, 76, 77, 106, 121, 134, 145, 159, 160, 175]. NSW [36, 37, 38, 39]. NSWB8 [39]. NSWTP [38]. NTIS [129]. Number [74]. Numbering [47]. Numbers [93, 117, 127].

October [63, 66, 160]. Operating [136]. Options [92, 114]. Overview [164].

Package [36]. Packet [8, 17, 78, 146]. Packet-switched [17]. Packets [163]. Part [184, 187, 188, 189]. Partitions [146]. PDP [136]. PDP-11 [136]. PDP-11/23/34/35/40 [136]. Peace [137]. Performance [157, 181]. Plan [47, 57, 107, 194]. Planning [200]. Plea [137]. Policy [207]. Port [83]. Pre [125]. Pre-emption [125]. Principles [155]. Private [174]. Problems [178]. Process [132]. Programs [159]. Project [8, 96, 168, 200, 151]. Proposal [27, 46, 62, 161, 195]. Proposed [26, 49, 119, 186]. Protocol [9, 28, 40, 41, 54, 55, 71, 72, 79, 80, 81, 85, 88, 90, 92, 94, 97, 111, 112, 113, 114, 118, 119, 123, 124, 128, 131, 133, 148, 149, 152, 158, 192, 193, 195, 197, 211, 38, 209]. Protocols [2, 12, 130, 140, 207]. Provision [35]. Provisional [57]. Public [17, 47]. Pup [138].

Queues [108].

Radio [78]. Radios [146]. Realization [153, 154]. Recommendation [47]. Reconfiguration [107]. Related [122]. Reliable [12, 193]. Report [151, 200]. Representation [39]. Request [79]. Research [200]. RESET [75]. Resources [62]. Retransmission [15, 50]. Revisions [18]. Rosen [191]. rough [9]. Routes [147]. Routing [19, 24, 25, 30, 46, 86, 95, 120, 144, 156, 162, 179, 181, 189, 190, 196]. Routings [23, 31].

Satellite [8]. SATNET [35, 42, 107, 192]. Scheme [49]. September [69, 121]. Sequence [74]. Sequences [150]. Server [61, 89, 103, 116, 142]. Service [17, 35, 51, 155, 201, 153, 154]. Services [185]. Short [201]. SIMP [9]. Simple [161, 169, 172]. Slow [147]. Some [52, 159, 181]. Source [95, 144]. Specification [4, 5, 7, 21, 28, 30, 40, 41, 54, 55, 148, 149, 186, 209, 211, 9]. SRI [136]. ST [119]. Standard [123, 124, 128, 129]. Standardization [152, 207]. Stanford [151]. Status [98, 155]. Stream [119]. Suggestion [180]. Summary [99, 101, 126]. supersedes [154]. Supervisor [4]. Support [161, 168]. switched [17]. Synchronization [173]. System [4, 136, 141, 164, 169]. Systems [178].

Table [94, 118]. TAC [166]. Tactical [110]. TCP [55, 151, 2, 5, 15, 18, 20, 21, 27, 29, 45, 49, 57, 64, 65, 66, 67, 68, 69, 70, 73, 75, 77, 98, 130, 150, 153, 154, 166, 167, 176]. TCP-4 [130]. TCP/IP [166, 176]. Techniques [15, 50]. Teleprocessing [84]. Telnet [148]. Term [201]. Terminal [132]. Test [79]. TFTP [133]. Thoughts [14, 15, 52]. Time [142, 173]. Timer [193]. Timer-Based [193]. Transaction [38]. Transfer [59, 149]. Transmission [17, 40, 55, 81, 112, 124, 129]. Transnet [35, 59]. Transport [144, 153, 154, 155, 162, 193, 195]. Types [51].

UCL [59]. UK [185, 190]. UK/US [185]. University [151, 185]. UNIX [168, 198]. Urgent [29]. USC [211]. USC/ISI [211]. User [13, 71, 88, 108, 176]. Users [161]. Utilizing [147].

VAX [168]. VAX-UNIX [168]. Verification [211]. Version [5, 27, 40, 41, 54, 55, 158]. Virtual [43]. Visit [171]. Voice [4].

Wars [137]. WBnet [165]. Wide [143, 144]. Work [211]. Working [60]. Worm [159].

X.25 [7, 130]. XNET [158].

Yellow [155, 153, 154].

References

IEN001

[1] C. J. Bennett, S. W. Edge, and A. J. Hinchley. Issues in the interconnection of datagram networks, July 29, 1977. URL http://www.cis.

     ohio-state.edu/htbin/ien/ien1.html. his document is also known as

INDRA note 637 and PSPWN 76.

IEN002

[2] Jon Postel. Comments on Internet protocols and TCP, August 15, 1977.

     URL http://www.cis.ohio-state.edu/htbin/ien/ien2.html.

IEN003

[3] Jon Postel. Internet meeting notes — 15 August 1977, August 18, 1977.

     URL http://www.cis.ohio-state.edu/htbin/ien/ien3.html.

IEN004

[4] ?. Rubin. Specification for a general network conferencing system voice communication supervisor, May 16, 1977. URL http://www.cis.ohio-

     state.edu/htbin/ien/ien4.html.

IEN005

[5] Vint Cerf. TCP version 2 specification, March 1977. URL http://www.

     cis.ohio-state.edu/htbin/ien/ien5.html.

IEN006

[6] Vint Cerf. Gateways and network interfaces, August 23, 1977. URL http:

     //www.cis.ohio-state.edu/htbin/ien/ien6.html.

IEN007

[7] Unknown. X.25 specification, July 1977. URL http://www.cis.ohio-

state.edu/htbin/ien/ien7.html.

IEN008

[8] UCL. Gateway experiments in connection with the packet satellite project,

???? ??, 19xx. URL http://www.cis.ohio-state.edu/htbin/ien/
ien8.html.

IEN009

[9] Radia Perlman. Host/SIMP protocol: A rough design specification,

August 26, 1977. URL http://www.cis.ohio-state.edu/htbin/ien/
ien9.html.

IEN010

[10] William W. Plummer. Internet broadcast issues, March 7, 1977. URL

http://www.cis.ohio-state.edu/htbin/ien/ien10.html.

IEN011

[11] Danny Cohen. Internetting or beyond NCP, March 21, 1977. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien11.html. This memo is also

assigned the numbers PRTN 213 and NSC 106.

IEN012

[12] L. Garlick, R. Rom, and J. Postel. Issues in reliable host-to-host pro-

tocols, March 15, 1977. URL http://www.cis.ohio-state.edu/htbin/
ien/ien12.html.

IEN013

[13] UCL. GNOME user’s guide, ???? ??, 19xx. URL http://www.cis.ohio-

state.edu/htbin/ien/ien13.html.

IEN014

[14] ?. Nelson. Thoughts on multi-net control and data collection facilities,

February 28, 1977. URL http://www.cis.ohio-state.edu/htbin/ien/
ien14.html.

IEN015

[15] William W. Plummer. Thoughts about TCP retransmission techniques,

September 15, 1977. URL http://www.cis.ohio-state.edu/htbin/
ien/ien15.html.

IEN016

[16] Jon Postel. Extensible field addressing, May 20, 1977. URL http://www.

cis.ohio-state.edu/htbin/ien/ien16.html.

IEN017

[17] ?. Bochman. Datagrams as a public packet-switched data transmission ser-

vice, March 1977. URL http://www.cis.ohio-state.edu/htbin/ien/
ien17.html.

IEN018

[18] ?. Burchfiel. TCP revisions, October 26, 1976. URL http://www.cis.

ohio-state.edu/htbin/ien/ien18.html.

IEN019

[19] John F. Shoch. Inter-network naming, addressing, and routing, January

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien19.html.

IEN020

[20] John F. Shoch. Inter-network fragmentation and the TCP, January 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien20.html.

IEN021

[21] Vint Cerf. TCP 3 specification, January 1978. URL http://www.cis.

ohio-state.edu/htbin/ien/ien21.html.

IEN022

[22] Jon Postel. Internet meeting notes — 1 February 1978, February 3, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien22.html.

IEN023

[23] Danny Cohen. On names, addresses and routings, January 23, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien23.html.

IEN024

[24] Radia Perlman. Gateway routing, January 27, 1978. URL http://www.

cis.ohio-state.edu/htbin/ien/ien24.html.

IEN025

[25] Virginia Strazisar. Gateway dynamic routing, January 25, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien25.html.

IEN026

[26] Vint Cerf. A proposed new Internet header format, February 14, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien26.html.

IEN027

[27] Vint Cerf. A proposal for TCP version 3.1 header format, February 14,

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien27.html.

IEN028

[28] Jon Postel. Draft Internetwork protocol specification, February 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien28.html.

IEN029

[29] ?. Day. A note on urgent in TCP, ???? ??, 19xx. URL http://www.cis.

ohio-state.edu/htbin/ien/ien29.html.

IEN030

[30] Virginia Strazisar. Gateway routing: An implementation specification,

April 11, 1978. URL http://www.cis.ohio-state.edu/htbin/ien/
ien30.html.

IEN031

[31] Dan Cohen. On names, addresses and routings (II), April 28, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien31.html.

IEN032

[32] John Davidson. Catenet monitoring and control: A model for the gate-

way component, April 28, 1978. URL http://www.cis.ohio-state.edu/
htbin/ien/ien32.html.

IEN033

[33] C. J. Bennett. Internet meeting notes — 1 & 2 May 1978, May 15, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien33.html.

IEN034

[34] C. J. Bennett. The GNOME controller, April 25, 1978. URL http://

www.cis.ohio-state.edu/htbin/ien/ien34.html.

IEN035

[35] P. T. Kirstein and C. J. Bennett. SATNET and the provision of transnet

service, April 28, 1978. URL http://www.cis.ohio-state.edu/htbin/
ien/ien35.html.

IEN036

[36] ?. Cashman. NSW file package, December 27, 1976. URL http://www.

cis.ohio-state.edu/htbin/ien/ien36.html.

IEN037

[37] ?. Millstein. NSW bibliography, ???? ??, 19xx. URL http://www.cis.

ohio-state.edu/htbin/ien/ien37.html.

IEN038

[38] Jon Postel. NSW Transaction Protocol (NSWTP), May 17, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien38.html.

IEN039

[39] Jon Postel. NSW data representation (NSWB8), May 17, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien39.html.

IEN040

[40] Jon Postel. Specification of Internetwork transmission control protocol —

version 4, June 1978. URL http://www.cis.ohio-state.edu/htbin/
ien/ien40.html.

IEN041

[41] Jon Postel. Internetwork protocol specification — version 4, June 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien41.html.

IEN042

[42] Winston Edmond. SATNET information, June 12, 1978. URL http://

www.cis.ohio-state.edu/htbin/ien/ien42.html.

IEN043

[43] ?. Pouzin. Interconnection of virtual circuits and datagrams, February

1977. URL http://www.cis.ohio-state.edu/htbin/ien/ien43.html.

IEN044

[44] Jon Postel. Latest header formats, June 1978. URL http://www.cis.

ohio-state.edu/htbin/ien/ien44.html.

IEN045

[45] William W. Plummer. TCP checksum function design, June 5, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien45.html.

IEN046

[46] David (?) Clark. A proposal for addressing and routing in the Internet,

June 1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien46.
html.

IEN047

[47] Unknown. CCITT draft recommendation on the numbering plan for public

data networks, ???? ??, 19xx. URL http://www.cis.ohio-state.edu/
htbin/ien/ien47.html.

IEN048

[48] Vint Cerf. The catenet model for Internetworking, July 1978. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien48.html.

IEN049

[49] P. L. Higginson. Comments on the proposed TCP checksum scheme, July

24, 1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien49.
html.

IEN050

[50] S. W. (?) Edge. A comparative evaluation of retransmission techniques,

???? ??, 19xx. URL http://www.cis.ohio-state.edu/htbin/ien/
ien50.html.

IEN051

[51] C. J. Bennett. Types of service on the catenet, July 21, 1978. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien51.html.

IEN052

[52] D. Cohen. Some thoughts about the multiplexing issue in networks,

August 11, 1978. URL http://www.cis.ohio-state.edu/htbin/ien/
ien52.html.

IEN053

[53] Jon Postel. Internet meeting notes — 2, 3 & 4 August 1978, August 21,

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien53.html.

IEN054

[54] Jon Postel. Internetwork protocol specification version 4, September 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien54.html.

IEN055

[55] Jon Postel. Specification of Internetwork Transmission Control Protocol TCP Version 4, September 1978. URL http://www.cis.ohio-state.

edu/htbin/ien/ien55.html.

IEN056

[56] David (?) Reed. CRC checksum calculation, September 11, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien56.html.

IEN057

[57] William W. Plummer. Provisional TCP development plan, October 5,

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien57.html.

IEN058

[58] Radia Perlman. Access control: An informal discussion, October 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien58.html.

IEN059

[59] C. J. Bennett. The UCL transnet file transfer implementation, October 23,

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien59.html.

IEN060

[60] John (?) Davidson. Boston area meeting of the Internet working group to discuss interactions with gateways, October 17, 1978. URL http://www.

cis.ohio-state.edu/htbin/ien/ien60.html.

IEN061

[61] Jon Postel. Internet name server, October 27, 1978. URL http://www.

cis.ohio-state.edu/htbin/ien/ien61.html.

IEN062

[62] John (?) Davidson. Enriched Internet addressing of ARPANET resources — an interim proposal, October 27, 1978. URL http://www.cis.ohio-

state.edu/htbin/ien/ien62.html.

IEN063

[63] Jon Postel. Internet meeting notes — 30 & 31 October 1978, November 14,

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien63.html.

IEN064

[64] C. Sunshine. TCP meeting notes — 12 march 1977, March 12, 1978. URL

http://www.cis.ohio-state.edu/htbin/ien/ien64.html.

IEN065

[65] Jon Postel. TCP meeting notes — 14 & 15 July 1977, August 5, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien65.html.

IEN066

[66] Jon Postel. TCP meeting notes — 13 & 14 October 1977, October 21,

1977. URL http://www.cis.ohio-state.edu/htbin/ien/ien66.html.

IEN067

[67] Jon Postel. TCP meeting notes — 30 & 31 January 1978, February 8,

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien67.html.

IEN068

[68] Jon Postel. TCP meeting notes — 15 & 16 June 1978, June 27, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien68.html.

IEN069

[69] Jon Postel. TCP meeting notes — 18 & 19 September 1978, October 9,

1978. URL http://www.cis.ohio-state.edu/htbin/ien/ien69.html.

IEN070

[70] Jon Postel. TCP meeting notes — 4 December 1978, December 15, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien70.html.

IEN071

[71] David (?) Reed. User datagram protocol, January 21, 1979. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien71.html.

IEN072

[72] Jonathan B. Postel. Multiplexing protocol, January 23, 1979. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien72.html.

IEN073

[73] Norm (?) Abramovitz. A TCP implementation issue, January 23, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien73.html.

IEN074

[74] William W. Plummer. Sequence number arithmetic, September 21, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien74.html.

IEN075

[75] William W. Plummer. The TCP RESET mechanism, September 21, 1978.

URL http://www.cis.ohio-state.edu/htbin/ien/ien75.html.

IEN076

[76] Jon Postel. Internet meeting notes — 25&26 January 1979, February 7,

1979. URL http://www.cis.ohio-state.edu/htbin/ien/ien76.html.

IEN077

[77] Jon Postel. TCP meeting notes — 29 January 1979, February 7, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien77.html.

IEN078

[78] John F. Shoch. Internetwork experiments with the Bay Area packet radio

network, February 1979. URL http://www.cis.ohio-state.edu/htbin/
ien/ien78.html.

IEN079

[79] ?. Lozano-Perez. A protocol test facility: Request for comments, January

1979. URL http://www.cis.ohio-state.edu/htbin/ien/ien79.html.

IEN080

[80] Jon Postel. Internet datagram protocol, February 1979. URL http://

www.cis.ohio-state.edu/htbin/ien/ien80.html.

IEN081

[81] Jon Postel. Transmission control protocol, February 1979. URL http://

www.cis.ohio-state.edu/htbin/ien/ien81.html.

IEN082

[82] Noel Chiappa, David Clark, and David Reed. LCS net address format,

February 12, 1979. URL http://www.cis.ohio-state.edu/htbin/ien/
ien82.html.

IEN083

[83] P. T. (?) Kirstein. Addressing through port expanders, March 23, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien83.html.

IEN084

[84] Danny (?) Cohen. A modern application of teleprocessing, April 1, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien84.html.

IEN085

[85] Jonathan B. Postel. Internet message protocol, March 1979. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien85.html.

IEN086

[86] Radia Perlman. Extended Internet routing, April 5, 1979. URL http://

www.cis.ohio-state.edu/htbin/ien/ien86.html.

IEN087

[87] Radia Perlman. Internet flow and congestion control, April 9, 1979. URL

http://www.cis.ohio-state.edu/htbin/ien/ien87.html.

IEN088

[88] Jonathan B. Postel. User datgram protocol, May 2, 1979. URL http://

www.cis.ohio-state.edu/htbin/ien/ien88.html.

IEN089

[89] Jon Postel. Internet name server, May 2, 1979. URL http://www.cis.

ohio-state.edu/htbin/ien/ien89.html.

IEN090

[90] Danny Cohen and Jon Postel. Multiplexing protocol, May 2, 1979. URL

http://www.cis.ohio-state.edu/htbin/ien/ien90.html.

IEN091

[91] Jonathan B. Postel. Address mappings, May 2, 1979. URL http://www.

cis.ohio-state.edu/htbin/ien/ien91.html.

IEN092

[92] Jon Postel. Protocol options, May 2, 1979. URL http://www.cis.ohio-

state.edu/htbin/ien/ien92.html.

IEN093

[93] J. Postel. Assigned numbers, May 3, 1979. URL http://www.cis.ohio-

state.edu/htbin/ien/ien93.html.

IEN094

[94] J. Postel. Internet protocol handbook table of contents, May 2, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien94.html.

IEN095

[95] Danny Cohen and Jon Postel. Source routing, May 2, 1979. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien95.html.

IEN096

[96] Anita P. Skelton, Steven F. Holmgren, and David C. Wood. The MITRE

cablenet project, May 2, 1979. URL http://www.cis.ohio-state.edu/
htbin/ien/ien96.html.

IEN097

[97] Steve Holmgren. Flexible datagram protocol, May 1979. URL http://

www.cis.ohio-state.edu/htbin/ien/ien97.html.

IEN098

[98] Jon Postel, Mike Wingfield, Norm Abramovitz, Bob Braden, Gary Grossman, Bill Plummer, Jim Mathis, Aage Stensby, and Dave Clark. TCP implementation status, May 3, 1979. URL http://www.cis.ohio-state.

edu/htbin/ien/ien98.html.

IEN099

[99] P. L. Higginson and C. J. Bennett. NI FTP: Summary and assess-

ment, May 3, 1979. URL http://www.cis.ohio-state.edu/htbin/ien/
ien99.html.

IEN100

[100] C. J. Bennett and P. L. Higginson. Comparison of the DIN FTP and the

NI FTP, May 3, 1979. URL http://www.cis.ohio-state.edu/htbin/
ien/ien100.html.

IEN101

[101] ?. Forsdick. AUTODIN FTP summary, May 4, 1979. URL http://www.

cis.ohio-state.edu/htbin/ien/ien101.html.

IEN102

[102] ?. Forsdick. Comparison of AUTODIN FTP with NI FTP, May 4, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien102.html.

IEN103

[103] J. R. Pickens, E. J. Feinler, and J. E. Mathis. An experimental network information center name server (NICNAME), May 4, 1979. URL http:/

/www.cis.ohio-state.edu/htbin/ien/ien103.html.

IEN104

[104] Virginia Strazisar. Minutes of the fault isolation meeting, March 12, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien104.html.

IEN105

[105] David Flood Page. ARPA catenet monitoring and control, May 25, 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien105.html.

IEN106

[106] Jon Postel. Internet meeting notes — 8, 9, 10, and 11 May 1979, May

17, 1979. URL http://www.cis.ohio-state.edu/htbin/ien/ien106.
html.

IEN107

[107] ?. Binder. SATNET reconfiguration plan, May 1979. URL http://www.

cis.ohio-state.edu/htbin/ien/ien107.html.

IEN108

[108] William W. Plummer. Internet user queues, June 19, 1979. URL http:/

/www.cis.ohio-state.edu/htbin/ien/ien108.html.

IEN109

[109] Virginia Strazisar. How to build a gateway, August 1979. URL http://

www.cis.ohio-state.edu/htbin/ien/ien109.html.

IEN110

[110] Vint Cerf. Internet addressing and naming in a tactical environment,

August 31, 1979. URL http://www.cis.ohio-state.edu/htbin/ien/
ien110.html.

IEN111

[111] Jon Postel. Internet protocol, August 1979. URL http://www.cis.ohio-

state.edu/htbin/ien/ien111.html.

IEN112

[112] Jon Postel. Transmission control protocol, August 1979. URL http://

www.cis.ohio-state.edu/htbin/ien/ien112.html.

IEN113

[113] Jonathan B. Postel. Internet message protocol, August 1979. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien113.html.

IEN114

[114] J. Postel. Protocol options, August 1979. URL http://www.cis.ohio-

state.edu/htbin/ien/ien114.html.

IEN115

[115] J. Postel. Address mappings, August 1979. URL http://www.cis.ohio-

state.edu/htbin/ien/ien115.html.

IEN116

[116] J. Postel. Name server, August 1979. URL http://www.cis.ohio-state.

edu/htbin/ien/ien116.html.

IEN117

[117] J. Postel. Assigned numbers, August 1979. URL http://www.cis.ohio-

state.edu/htbin/ien/ien117.html.

IEN118

[118] J. Postel. Internet protocol handbook table of contents, August 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien118.html.

IEN119

[119] James W. Forgie. ST — A proposed Internet stream protocol, September

7, 1979. URL http://www.cis.ohio-state.edu/htbin/ien/ien119.
html.

IEN120

[120] Radia Perlman. Internet routing and the network, October 1979. URL

http://www.cis.ohio-state.edu/htbin/ien/ien120.html.

IEN121

[121] J. Postel. Internet meeting notes — 10, 11, 12, & 13, September 1979,

October 25, 1979. URL http://www.cis.ohio-state.edu/htbin/ien/
ien121.html.

IEN122

[122] Danny Cohen. On addressing and related issues (or: Fuel for a discussion),

October 31, 1979. URL http://www.cis.ohio-state.edu/htbin/ien/
ien122.html.

IEN123

[123] Jon Postel. DOD standard Internet protocol, December 1979. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien123.html.

IEN124

[124] Jon Postel. DOD standard transmission control protocol, December 1979.

URL http://www.cis.ohio-state.edu/htbin/ien/ien124.html.

IEN125

[125] V. Cerf. Pre-emption, December 1979. URL http://www.cis.ohio-

state.edu/htbin/ien/ien125.html.

IEN126

[126] Danny Cohen. Summary of the ARPA/ETHERNET community meet-

ing, November 1979. URL http://www.cis.ohio-state.edu/htbin/
ien/ien126.html.

IEN127

[127] J. Postel. Assigned numbers, January 1980. URL http://www.cis.ohio-

state.edu/htbin/ien/ien127.html.

IEN128

[128] Jon Postel. DOD standard Internet protocol, January 1980. URL http:

//www.cis.ohio-state.edu/htbin/ien/ien128.html.

IEN129

[129] Jon Postel. DOD standard transmission control (NTIS no. ADA079730),

January 1980. URL http://www.cis.ohio-state.edu/htbin/ien/
ien129.html.

IEN130

[130] Unknown. Comparison of X.25 and TCP-4 as cable-bus network proto-

cols, ???? ??, 19xx. URL http://www.cis.ohio-state.edu/htbin/ien/
ien130.html.

IEN131

[131] David Flood Page. Gateway monitoring protocol, February 1, 1980. URL

http://www.cis.ohio-state.edu/htbin/ien/ien131.html.

IEN132

[132] David Flood Page. The CMCC terminal process, February 1, 1980. URL

http://www.cis.ohio-state.edu/htbin/ien/ien132.html.

IEN133

[133] Karen R. Sollins. The TFTP protocol, January 29, 1980. URL http://

www.cis.ohio-state.edu/htbin/ien/ien133.html.

IEN134

[134] J. Postel. Internet meeting notes — 4, 5, & 6 February 1980, February

29, 1980. URL http://www.cis.ohio-state.edu/htbin/ien/ien134.
html.

IEN135

[135] C. Sunshine and J. Postel. Addressing mobile hosts in the ARPA Inter-

net environment, March 1980. URL http://www.cis.ohio-state.edu/
htbin/ien/ien135.html.

IEN136

[136] S. R. Wiseman and B. H. Davies. Memory management extensions to the SRI micro operating system for PDP-11/23/34/35/40, May 1, 1980. URL

http://www.cis.ohio-state.edu/htbin/ien/ien136.html.

IEN137

[137] Danny Cohen. On holy wars and a plea for peace, April 1, 1980. URL

http://www.cis.ohio-state.edu/htbin/ien/ien137.html. This note

discusses the Bit-Endian/Little-Endian byte/bit-order controversy, but did not settle it. A decade later, David V. James in “Multiplexed Buses: The Endian Wars Continue”, IEEE Micro, 10(3), 9–21 (1990) continued the discussion.

IEN138

[138] John F. Shoch. Initial comparison of EPRs and IPRs in the Pup Inter-

net environment, April 1980. URL http://www.cis.ohio-state.edu/
htbin/ien/ien138.html.

IEN139

[139] Jack Haverty. HOSTs as IMPs, April 1980. URL http://www.cis.ohio-

state.edu/htbin/ien/ien139.html.

IEN140

[140] John F. Shoch. Mutual encapsulation of Internetwork protocols, April

1980. URL http://www.cis.ohio-state.edu/htbin/ien/ien140.html.

IEN141

[141] C. J. Bennett. Message system issues, April 11, 1980. URL http://www.

cis.ohio-state.edu/htbin/ien/ien141.html.

IEN142

[142] K. Harrenstien and J. Postel. Time server, April 1980. URL http://www.

cis.ohio-state.edu/htbin/ien/ien142.html.

IEN143

[143] Jerome H. Saltzer. Environment considerations for campus-wide net-

works, March 1980. URL http://www.cis.ohio-state.edu/htbin/ien/
ien143.html.

IEN144

[144] Jerome H. Saltzer. Source routing for campus-wide Internet trans-

port, March 1980. URL http://www.cis.ohio-state.edu/htbin/ien/
ien144.html.

IEN145

[145] J. Postel. Internet meeting notes — 14 & 15 May 1980, May 29, 1980.

URL http://www.cis.ohio-state.edu/htbin/ien/ien145.html.

IEN146

[146] Radia Perlman. Flying packet radios and network partitions, June 1980.

URL http://www.cis.ohio-state.edu/htbin/ien/ien146.html.

IEN147

[147] Radia Perlman. Utilizing Internet routes as expressways through slow

nets, June 1980. URL http://www.cis.ohio-state.edu/htbin/ien/
ien147.html.

IEN148

[148] J. Postel. Telnet protocol specification, June 1980. URL http://www.

cis.ohio-state.edu/htbin/ien/ien148.html.

IEN149

[149] J. Postel. File transfer protocol specification, June 1980. URL http://

www.cis.ohio-state.edu/htbin/ien/ien149.html.

IEN150

[150] W. Plummer and J. Postel. TCP JSYS calling sequences, June 1980. URL

http://www.cis.ohio-state.edu/htbin/ien/ien150.html.

IEN151

[151] Vint Cerf. Final report of the Stanford University TCP Project, April 1,

1980. URL http://www.cis.ohio-state.edu/htbin/ien/ien151.html.

IEN152

[152] Vint Cerf. DoD protocol standardization, July 1, 1980. URL http://

www.cis.ohio-state.edu/htbin/ien/ien152.html.

IEN153

[153] C. J. Bennett. Realization of the Yellow Book Transport Service above

TCP, July 29, 1980. URL http://www.cis.ohio-state.edu/htbin/
ien/ien153.html.

IEN154

[154] C. J. Bennett. Realization of the Yellow Book Transport Service above TCP (supersedes IEN 153), July 29, 1980. URL http://www.cis.ohio-

state.edu/htbin/ien/ien154.html.

IEN155

[155] C. J. Bennett. The yellow book transport service: Principles and status,

August 12, 1980. URL http://www.cis.ohio-state.edu/htbin/ien/
ien155.html.

IEN156

[156] Danny Cohen. Controlled routing in the catenet, September 7, 1980. URL

http://www.cis.ohio-state.edu/htbin/ien/ien156.html.

IEN157

[157] David Flood Page. CMCC performance measurement message formats,

September 21, 1980. URL http://www.cis.ohio-state.edu/htbin/
ien/ien157.html.

IEN158

[158] Jack Haverty. XNET formats for Internet protocol version 4, October 1,

1980. URL http://www.cis.ohio-state.edu/htbin/ien/ien158.html.

IEN159

[159] John F. Shoch and Jon A. Hupp. Notes on the “worm” programs — some early experience with a distributed computation, October 1980. URL

http://www.cis.ohio-state.edu/htbin/ien/ien159.html.

IEN160

[160] J. Postel. Internet meeting notes — 7-8-9 October 1980, November 7, 1980.

URL http://www.cis.ohio-state.edu/htbin/ien/ien160.html.

IEN161

[161] R. G. Jones. A proposal for simple measurement support for users, Novem-

ber 1980. URL http://www.cis.ohio-state.edu/htbin/ien/ien161.
html.

IEN162

[162] John A. Pershing, Jr. Transport, addressing, and routing in, October 1980.

URL http://www.cis.ohio-state.edu/htbin/ien/ien162.html.

IEN163

[163] R. G. Jones. Echo delay measurements with GGP packets, November

1980. URL http://www.cis.ohio-state.edu/htbin/ien/ien163.html.

IEN164

[164] Jerry Stern, Greg Ruth, and Jack Haverty. CMOS system overview, Jan-

uary 1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien164.
html.

IEN165

[165] Danny Cohen, Jon Postel, and Steve Casner. About addressing in the

WBnet, January 1981. URL http://www.cis.ohio-state.edu/htbin/
ien/ien165.html.

IEN166

[166] Robert Hinden. Design of TCP/IP for the TAC, January 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien166.html.

IEN167

[167] Jack Sax and Winston Edmond. HP3000 TCP design document, July

1980. URL http://www.cis.ohio-state.edu/htbin/ien/ien167.html.

IEN168

[168] Robert F. Gurwitz. VAX-UNIX networking support project implementation description, January 1981. URL http://www.cis.ohio-state.edu/

htbin/ien/ien168.html.

IEN169

[169] C. J. Bennett. A simple NIFTP-based mail system, January 23, 1981.

URL http://www.cis.ohio-state.edu/htbin/ien/ien169.html.

IEN170

[170] Danny Cohen and Jon Postel. On IP-addressing, January 25, 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien170.html.

IEN171

[171] Danny Cohen. Addressing in the ARPAnet, another visit, January 18,

1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien171.html.

IEN172

[172] David Flood Page. A simple message generator, March 9, 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien172.html.

IEN173

[173] D. L. Mills. Time synchronization in DCNET hosts, February 25, 1981.

URL http://www.cis.ohio-state.edu/htbin/ien/ien173.html.

IEN174

[174] International Telegraph and Telephone Consultative Committee (CCITT). Private network DTE addressing, February 12, 1981. URL http://www.

cis.ohio-state.edu/htbin/ien/ien174.html.

IEN175

[175] J. Postel. Internet meeting notes — 28-29-30 January 1981, March 13,

1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien175.html.

IEN176

[176] Dave Lyons. The DECSYSTEM-20 TCP/IP user interface, March 1981.

URL http://www.cis.ohio-state.edu/htbin/ien/ien176.html.

IEN177

[177] J. Postel. Comments on action items from the January meeting, March

24, 1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien177.
html.

IEN178

[178] Carl A. Sunshine. Addressing problems in multi-network systems, April

1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien178.html.

IEN179

[179] Danny Cohen. Addressing and routing, March 11, 1981. URL http://

www.cis.ohio-state.edu/htbin/ien/ien179.html.

IEN180

[180] Danny Cohen. A suggestion for Internet message forwarding for MO-

SIS, March 1981. URL http://www.cis.ohio-state.edu/htbin/ien/
ien180.html.

IEN181

[181] Jack Haverty. Van gateway: Some routing and performance issues, May

1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien181.html.

IEN182

[182] Eric C. Rosen. Issues in buffer management, May 1981. URL http://

www.cis.ohio-state.edu/htbin/ien/ien182.html.

IEN183

[183] Eric C. Rosen. Logical addressing, May 1981. URL http://www.cis.

ohio-state.edu/htbin/ien/ien183.html.

IEN184

[184] Eric C. Rosen. Issues in Internetting part 1: Modelling the Internet, May

1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien184.html.

IEN185

[185] Robert T. Braden and Peter L. Higginson. Development of UK/US network services at University College, London, May 1981. URL http://

www.cis.ohio-state.edu/htbin/ien/ien185.html.

IEN186

[186] Mary M. Bernstein. Proposed DCEC IP specification, May 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien186.html.

IEN187

[187] Eric C. Rosen. Issues in Internetting part 2: Accessing the Internet, June

1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien187.html.

IEN188

[188] Eric C. Rosen. Issues in Internetting part 3: Addressing, June 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien188.html.

IEN189

[189] Eric C. Rosen. Issues in Internetting part 4: Routing, June 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien189.html.

IEN190

[190] Robert Cole and Robert Braden. Routing and access control in UK to

US, July 9, 1981. URL http://www.cis.ohio-state.edu/htbin/ien/
ien190.html.

IEN191

[191] C. Sunshine, D. Cohen, and J. Postel. Comments on Rosen’s memos,

July 1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien191.
html.

IEN192

[192] Bolt Beranek and Newman Inc. Host/Satnet protocol, July 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien192.html.

IEN193

[193] Dick Watson. Timer-based mechanisms in reliable transport protocol connection management, July 1981. URL http://www.cis.ohio-state.

edu/htbin/ien/ien193.html.

IEN194

[194] D. L. Mills and M. J. O’Connor. DCNET mail plan, July 23, 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien194.html.

IEN195

[195] Carl Sunshine. Comments on NBS transport protocol proposal, August

1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien195.html.

IEN196

[196] B. Bowman and P. Cook. Issues involving non-routing gateways, Septem-

ber 1981. URL http://www.cis.ohio-state.edu/htbin/ien/ien196.
html.

IEN197

[197] Benjamin M. Littauer, Andrew J. Huang, and Robert M. Hinden. A host monitoring protocol, September 1981. URL http://www.cis.ohio-

state.edu/htbin/ien/ien197.html.

IEN198

[198] Robert Cole. Extended memory MOS for a Unix development host,

September 1981. URL http://www.cis.ohio-state.edu/htbin/ien/
ien198.html.

IEN199

[199] Unknown. Never issued, 19xx. URL http://www.cis.ohio-state.edu/

htbin/ien/ien199.html.

IEN200

[200] David D. Clark. Internet project research planning report, December 1981.

URL http://www.cis.ohio-state.edu/htbin/ien/ien200.html.

IEN201

[201] David D. Clark. Internet short term service goals, December 1981. URL

http://www.cis.ohio-state.edu/htbin/ien/ien201.html.

IEN202

[202] Never issued. Not yet issued, ???? ??, 19xx. URL http://www.cis.ohio-

state.edu/htbin/ien/ien202.html.

IEN203

[203] Never issued. Not yet issued, ???? ??, 19xx. URL http://www.cis.ohio-

state.edu/htbin/ien/ien203.html.

IEN204

[204] Never issued. Not yet issued, ???? ??, 19xx. URL http://www.cis.ohio-

state.edu/htbin/ien/ien204.html.

IEN205

[205] Never issued. Not yet issued, ???? ??, 19xx. URL http://www.cis.ohio-

state.edu/htbin/ien/ien205.html.

IEN206

[206] Never issued. Not yet issued, ???? ??, 19xx. URL http://www.cis.ohio-

state.edu/htbin/ien/ien206.html.

IEN207

[207] Richard D. DeLauer. Dod policy on standardization of host-to-host protocols for data communications networks, March 1982. URL http://www.

cis.ohio-state.edu/htbin/ien/ien207.html.

IEN208

[208] R. Hinden and A. Sheltzer. The DARPA Internet gateway, March 1982.

URL http://www.cis.ohio-state.edu/htbin/ien/ien207.html. This

note was changed to RFC 823 instead, and never issued as an IEN.

IEN209

[209] Dave Mills. Exterior Gateway Protocol (EGP) formal specification, March

1982. URL http://www.cis.ohio-state.edu/htbin/ien/ien207.html.

This note was changed to RFC 904 instead, and never issued as an IEN.

IEN210

[210] Never issued. Not yet issued, ???? ??, 19xx. URL http://www.cis.ohio-

state.edu/htbin/ien/ien210.html.

IEN211

[211] Carl Sunshine. Protocol specification and verification work at USC/

ISI, August 1982. URL http://www.cis.ohio-state.edu/htbin/ien/
ien211.html.

IEN212

[212] Robert Gurwitz and Robert Hinden. IP-local area network addressing

     issues, September 1982. URL http://www.cis.ohio-state.edu/htbin/
     ien/ien212.html.